OpenStudy (anonymous):

Solve the separable differential equation 11 x - 6 y √(x²+ 1) (dy/dx) = 0. Subject to the initial condition: y(0) = 4. y =

OpenStudy (unklerhaukus):

rearrange to the form f(y) dy = g(x) dx

OpenStudy (unklerhaukus):

can you do that @eescobedo1 ?

OpenStudy (unklerhaukus):

good , now integrate both sides ∫-6ydy =∫ -11/√(x²+1) (dx)

OpenStudy (anonymous):

ok what I got was \[-3y ^{2}=-11\sqrt{x^{2}+1} + c\]

OpenStudy (unklerhaukus):

now solve it for y(x),

OpenStudy (anonymous):

do you mean to use the initial condition at this point where I would let x=0 and y=4 to solve for c then substitute it back to the equation to solve for y?

OpenStudy (unklerhaukus):

yeah but i usually get the equation into y(x) = ... form before subbing in initlal condition

OpenStudy (unklerhaukus):

for this equation though it might be more sensible to just get it into y^2(x) = . . ., because you dont want to forget that extra solution

OpenStudy (anonymous):

ok well then this is where I'm getting lost....

OpenStudy (unklerhaukus):

what do you get for y^2

OpenStudy (anonymous):

what I'm thinking is to divide the -3 by both sides to get \[y ^{2}=11\sqrt{x ^{2}+1}+c/3\]

OpenStudy (unklerhaukus):

yeah so the result is \[y^2(x)=\frac{11\sqrt{x^2+1}+c}3\] if we change the constant of integration a bit \(C=c/3\) \[y^2(x)=\tfrac{11}3\sqrt{x^2+1}+C\]

OpenStudy (unklerhaukus):

Now we can sub in y(0) = 4, ie x=0, y=4

OpenStudy (unklerhaukus):

and solve for C (or c)

OpenStudy (anonymous):

ok let me try =)

OpenStudy (unklerhaukus):

actually i see now that \(C\) is not a good idea, it is simpler calculation with \(c\), but they will give the same equation eventually

OpenStudy (anonymous):

ok so I got c=-11, from this point I would sub my answer to the equation: \[y ^{2}=11/3\sqrt{x ^{2}+1}-11\]

OpenStudy (unklerhaukus):

that would be right, but i think there is a calculation error

OpenStudy (unklerhaukus):

\[16 \neq \tfrac{11}3 - 11\]

OpenStudy (anonymous):

ok because the first time I worked it out I got c=-37 so at this point I would have \[y ^{2}=11/3\sqrt{x ^{2}+1}+37/3\]

OpenStudy (unklerhaukus):

yeah that's right \[y ^{2}(x)=\frac{11\sqrt{x ^{2}+1}+37}3\]

OpenStudy (unklerhaukus):

and we can check the initial point works 4^2 = (11+37) /3 = 16 it does!

OpenStudy (unklerhaukus):

if you really wanted to, you could get \[y(x) = \pm\sqrt{\dots}\] but \[y^2(x) = \dots \] is much nicer

OpenStudy (anonymous):

I know because we have to square both sides in or to get y, which is what they are asking, it just looks messy being that there is a square root over another square root...

OpenStudy (unklerhaukus):

yep, y(x) does look messy, even if you write the sqrts as ^{1/2}, not much you can do about it, that why y^2(x)= ... is nicer

OpenStudy (unklerhaukus):

or maybe even \[3y^2(x) = 11\sqrt{x^2+1}+ 37\] is nicier still

OpenStudy (anonymous):

Thank you for your help, pretty much I entered the answer as \[\pm \sqrt{....}\] and received an incorrect response of: Your answer isn't a formula that returns a number (it looks like a formula that returns a list of numbers

OpenStudy (unklerhaukus):

hmm, i would try entering y^2(x) = . . .

OpenStudy (anonymous):

Ok after many attempts to input the answer correctly I finally was able to input it in the correct form and yes our answer was correct! I was just having trouble entering it in the program I am using, much thanks with helping me!!!

OpenStudy (unklerhaukus):

oh wait if you check the detail \[y ^{2}(x)=\frac{11\sqrt{x ^{2}+1}+37}3\\ y(x) = \pm\sqrt{\frac{11\sqrt{x ^{2}+1}+37}3}\] remember \[y(0) =4 \] the negative solution dosen't fit the initial condition

OpenStudy (unklerhaukus):

because the sqrt term must be positive,

OpenStudy (anonymous):

and you are correct I didn't submit the negative answer! =)

OpenStudy (unklerhaukus):

the plus/minus should only be there before we find c

OpenStudy (unklerhaukus):

ok, well i have to go, Welcome to OpenStudy!